Wilmington Company has two manufacturing departments--Assembly and Fabrication. It considers all of its manufacturing overhead costs to be fixed costs. The first set of data that is shown below is based on estimates from the beginning of the year. The second set of data relates to one particular job completed during the year--Job Bravo.

Estimated Data Assembly Fabrication Total
Manufacturing overhead costs $ 1,650,000 $ 1,950,000 $ 3,600,000
Direct labor-hours 75,000 45,000 120,000
Machine-hours 30,000 150,000 180,000
Job Bravo Assembly Fabrication Total
Direct labor-hours 16 8 24
Machine-hours 8 11 19

Required:

1. If Wilmington used a plantwide predetermined overhead rate based on direct labor-hours, how much manufacturing overhead would be applied to Job Bravo?

2. If Wilmington uses departmental predetermined overhead rates with direct labor-hours as the allocation base in Assembly and machine-hours as the allocation base in Fabrication, how much manufacturing overhead would be applied to Job Bravo?

(Round your intermediate calculation to 2 decimal places.)
Answer #1

If Wilmington used a plantwide predetermined overhead rate based on direct labor-hours, how much manufacturing overhead would be applied to Job Bravo?

Plantwide predetermined overhead rate=Estimated Manufacturing overhead costs/Total Direct labor-hours

=3,600,000/120,000=$30 per direct labor hr

how much manufacturing overhead would be applied to Job Bravo?

=24 hr *$30 per direct labor =$720

2. If Wilmington uses departmental predetermined overhead rates with direct labor-hours as the allocation base in Assembly and machine-hours as the allocation base in Fabrication, how much manufacturing overhead would be applied to Job Bravo?

Assembly overhead rate =1,650,000/75,000=$22 per direct labor hr

Fabrication overhead rate =1,950,000/150000=$13 per Machne hr

how much manufacturing overhead would be applied to Job Bravo?

Assembly=16*$22 per direct labor hr=$352

Fabrication=11*$13 per Machne hr=$143

Total Manufacturing overhead applied =$352+$143=$495